### v2

# Evolving a Multi-Population Evolutionary-QAOA on Distributed QPUs

Abstract
Our research combines an Evolutionary Algorithm (EA) with a Quantum Approximate Optimization Algorithm (QAOA) to update the ansatz parameters, in place of traditional gradient-based methods, and benchmark on the Max-Cut problem. We demonstrate that our Evolutionary-QAOA (E-QAOA) pairing performs on par or better than a COBYLA-based QAOA in terms of solution accuracy and variance, for d-3 regular graphs between 4 and 26 nodes, using both max_count and Conditional Value at Risk (CVaR) for fitness function evaluations. Furthermore, we take our algorithm one step further and present a novel approach by presenting a multi-population EA distributed on two QPUs, which evolves independent and isolated populations in parallel, classically communicating elite individuals. Experiments were conducted on both simulators and IBM quantum hardware, and we investigated the relative performance accuracy and variance.
Index Terms:
Evolutionary Algorithm, QAOA, Optimisation, Max-Cut, Distributed Computing, Multi-population Algorithm
I Introduction
Quantum Approximate Optimization Algorithms (QAOAs) are promising methods for solving combinatorial optimization problems [1, 2, 3] on quantum computers. The original QAOA presented by Farhi, Goldstone and Gutmann in 2014 [4] demonstrated the potential for quantum advantage as it outperformed the best-known classical algorithm to solve the Max-Cut problem at the time. Since then classical computing has taken greater strides to maintain an advantage over quantum optimization. Yet, the QAOA continues to be an active area of research for the quantum computing community [5]. The QAOA is a variational algorithm [6] and therefore faces the well-known challenges of barren plateaus and noise degradation [7, 8] that occur with increasing circuit width and depth in the current era of noisy, near-term quantum computers. These issues make it increasingly hard for optimizers to move in a search landscape as the problem grows to a relevant scale [9, 10]. These difficulties motivate our research into an Evolutionary Algorithm (EA) as an optimizer for QAOA, instead of linear or gradient-based methods. There is still a discussion within the literature about whether or not gradient-free methods are useful, for example, Arrasmith et al. argue that these do not solve the barren plateau problem [11]. However, others state that Genetic Algorithms (GAs) and EAs have an advantage for faster convergence [12], avoiding barren plateaus [13], and beating the state-of-the-art gradient-free optimization for better accuracy [14]. Note that there is no clear consensus in the field of the distinction between GA and EA and the two have overlapping definitions [15]. As such, a benefit for one would clearly benefit the other. Furthermore, EAs have been shown to handle complex optimization problems with noisy or uncertain fitness functions [15, 16]. If these conditions hold, and EAs prove to be more robust as the literature continues to tackle larger-scale problems, this would make them a natural choice in combining them with QAOAs and other Variational Quantum Algorithms (VQAs).
In addition, EAs offer the potential for embarrassingly parallel applications, one specific method developed is that of Multi-population Genetic Algorithms (MGAs) [17, 18], which evolve independent populations in isolation. MGAs have shown to be successful in avoiding premature convergence, an issue that arises sometimes in GAs, and increasing diversity within the population’s evolution, a key factor for successful convergence [17, 19, 20]. We present a technique that implements an MGA, combining a multi-population evolutionary algorithm with a QAOA. We propose to distribute independent population instances on two separate Quantum Processing Units (QPUs) orchestrated by a scheduler. This strategy is novel to the best of the authors’ knowledge: numerous scheduling prescriptions have been proposed by the community [21, 22, 23], however, none are tailored to parallelizing MGAs across QPUs.
This paper reports on the solution accuracy and variance of our EA compared to the traditional Constrained Optimization BY Linear Approximation (COBYLA) method [24, 25, 26] used with the QAOA. Tests were performed on regular $d$-3 graphs both on the simulator and real quantum hardware. The paper is structured as follows. The methodology in Sec. II provides a focused overview of the QAOA for a Max-Cut problem and the relevant elements for the EA. Sec. III outlines the computational facilities used and the parameters chosen in our experiments. Next, we show the results in Sec. IV, together with their discussion. Lastly, Sec. V summarizes the work presented and its limitations and concludes with an outlook on future work.
II Methodology
In this study, we present the components of our hybrid evolutionary QAOA, referred to as E-QAOA. Firstly, we describe the problem we are trying to solve, the Max-Cut problem. Secondly, we detail how to construct the quantum circuit and how to process measurement outcomes. Thirdly, we give an overview of the EA elements required to evolve a population of solutions. Note that this section will also give a brief overview of the terminology used in the fields of evolutionary and genetic algorithms, to then use them throughout the rest of the paper. It is worth mentioning that in the EA literature, the functional cost is also referred to as fitness; as such these two terms will be used interchangeably.
II-A Max-Cut
As mentioned already, the first QAOA published in the literature [4] used a variational quantum circuit to solve the Max-Cut problem. Where the solution to this problem can be neatly summarized as stating that given an unweighted graph, with $n$ nodes, and $e$ edges, the Max-Cut of this graph is given by the boundary which divides the nodes into two sets, such that the number of edges between the two sets is maximized. In the quantum setting, there is a one-to-one mapping between nodes and qubits. The quantum algorithm samples bit-strings $x\in\{0,1\}^{n}$ where the set that a given node belongs to is labelled as ‘0’ or ‘1’.
II-B The QAOA circuit
Figure 1: Example circuits for $d$-3 regular graphs with 4 (top) and 6 (bottom) nodes. The circuits have as many qubits as there are in the respective graphs, and are structured as follows: a Hadamard gate ($H$, in red) is first applied to all qubits, followed by $RZZ$ gates constituting the connectivity section and $Rx$ gates on all qubits. This circuit section (dotted-line rectangle) is repeated $p$ times, before the qubit states are measured. Crucially, increasing either $p$ or the number of edges, i.e. the graph connectivity, increases the circuit depth.
We chose a basic QAOA without enhancements, to effectively study the effects of pairing an EA to it. The version we chose is described in Qiskit’s documentation [27]. In the plain algorithm, the circuit is composed of parameterised time evolution unitaries that are alternately generated by a cost Hamiltonian ($H_{C}$) and a mixer Hamiltonian ($H_{M}$). Given the Max-Cut problem, as described above, the unitaries of these are given by
$U(H_{M})$ $=e^{-i\beta H_{M}}=e^{-i\beta X_{0}}e^{-i\beta X_{1}}\cdots e^{-i\beta X_{n}}$ (1)
$U(H_{C})$ $=e^{-i\gamma H_{C}}=e^{-i\gamma Z_{i}Z_{j}}e^{-i\gamma Z_{j}Z_{k}}\cdots e^{-i\gamma Z_{k}Z_{i}},$ (2)
where $X_{i}$ and $Z_{i}$ are the Pauli X and Z matrices respectively, acting on the $i^{\rm th}$ qubit (node). The $RZZ$ gates are applied to qubit pairs $(i,j)$ that correspond to edges in the graph. The angle of these rotations is determined by the ansatz parameters $\beta$ and $\gamma$. The unitary $U(H_{M})$ and $U(H_{C})$ blocks are repeated $p$ times, where $p$ is also known as the number of rounds in QAOA. As $\lim_{p\to\infty}$, the algorithm converges to the optimal solution, given the correct optimization of the ansatz parameters. The parameters $\beta$ and $\gamma$ change for each $p$ block, and are initialized randomly on $(-\pi,\pi]$. The resulting circuit structure for a given graph is aided by a simple diagram in Fig. 1. While many optimizers can be chosen for QAOA, the traditional choice in the literature and documentation used to update $\beta$ and $\gamma$ is COBYLA, which is a linear-based optimiser, and in this work, it was used through the minimize method in Scipy [28].
II-C Fitness evaluation
In our experiments, we considered two metrics to measure the fitness of the solution, and thus update $\beta$ and $\gamma$. In the first instance, the fitness was defined as the Max-Cut number of the most frequent bit-string, rather than the average or expected output, given by the measured distribution of each circuit simulation. This can be obtained through Qiskit’s max_count feature. In the second instance, we used Conditional Value-at-Risk (CVaR), which takes the expectation value of the best $\alpha$ partitions found in the output distribution [29]. Where $\alpha$ is a value between 0 and 1, and $\alpha=1$ is equivalent to the standard expectation value using the whole distribution. We chose this method as CVaR leads to faster convergence on both simulators and hardware, as tested on various combinatorial optimization problems [30]. Finally, rather than using the Max-Cut value of these partitions directly, we use the approximation ratio, which is simply the Max-Cut value we find divided by the true optimal Max-Cut solution, such that this value is normalized on 1 for all graph sizes, and we can see how close we are to the true solution:
[…]
II-D The Evolutionary algorithm
In our method, we replace COBYLA with an EA that is used to update the variational parameters. All evolutionary or genetic algorithms have some core components to their structure and may slightly change in things like mutation operations or reproduction techniques, but the field allows the two to be used interchangeably. The core components describing an EA are thus: parent selection, recombination operator, mutation, and survivor selection. These choices of strategy help maintain uniqueness and diversity within the population genome such that the algorithm can achieve a good balance between exploitation and exploration in its evolution and research of the fitness landscape. There are other factors as well that can compose an evolutionary algorithm, like population size, replacement strategy, termination strategy and more, but these are less important than the core components. We will limit our explanation of the background theory of this field to the components that we made use of, and we will lead the readers through the terminology that we mention. In an EA we start with a population of solutions, so rather than just one initial random point for most optimization methods, we start with $n_{\rm pop}$ values. These are also initialized randomly, and each individual in the population, denoted by $Y=\{Y_{1},Y_{2},\cdots,Y_{n_{\rm pop}}\}$, is described the same as in COBYLA, with a real-valued array between $-\pi$ and $\pi$ of the form
$Y_{i}=[\beta_{1},\gamma_{1},\beta_{2},\gamma_{2},\cdots,\beta_{p},\gamma_{p}]\,.$ (4)
In evolutionary terminology, this array is defined as the genotype or chromosome, and all of the values within this array are referred to as genes or alleles. As the reader will note, the terminology within the field adopts terms from biological evolution theory. In the metaphor that compares Darwinian evolution to a programmed evolution, GAs use these notions which make it easier for the reader and users of it to understand the concepts being used, how they form the algorithm, and why they should operate. Thus, the value of $Y$ is described as the genotype because in evolution the genetic code of an individual will be the factor determining its fitness relative to the population, its likelihood to survive and reproduce and carry forth its genetic code in future generations. From this, the algorithm follows. Once we have our population, we determine the fitness of each $Y_{i}$ and use that value to choose the individuals that will pair up as parents and reproduce to form a new generation of solutions or offspring. The parent selection strategy is given by Stochastic Universal Sampling (SUS) [31] where each individual’s probability of being chosen as a parent is proportional to their normalized fitness value. This ensures that all individuals have a chance to carry their genes over to the next generation, apart from the weakest link. We have chosen for the population size to remain constant, such that two parents produce two new offspring until a population of the same size is achieved. SUS allows sampling with replacement, where the fittest individual can be chosen in multiple, or all pairs of parents, but can only be selected once for each pair of parents and thus they cannot pair with themselves. Once the parents are chosen, the offspring are produced via recombination and mutation. Recombination was applied $100\%$ of the time, on each pair of parents, through whole arithmetic crossover, i.e. on all of their genes. This operation is described below, where, given two parents, $Y_{i}$ and $Y_{j}$, one of their offspring $Y_{i}^{\prime}$ is given by
$\begin{split}Y^{\prime}_{i}=[~{}&\alpha\beta_{Y_{i},1}+(1-\alpha)\beta_{Y_{j},1},\alpha\gamma_{Y_{i},1}+(1-\alpha)\gamma_{Y_{j},1},\\ &\cdots,\\ &\alpha\beta_{Y_{i},p}+(1-\alpha)\beta_{Y_{j},p},\alpha\gamma_{Y_{i},p}+(1-\alpha)\gamma_{Y_{j},p}~{}],\end{split}$ (5)
where $\alpha$ is a randomly generated variable uniform on $(0,1]$. Mutation is then applied to each of the genes with a probability $p_{\sigma}=0.2$, this again follows the concept of natural evolution, where offspring are never the exact combination of their parents, but in fact, undergo random mutation on their genetic codes. In our EA we use self-adaptive mutation, where we add $2p$ different $\sigma$ mutation values within the genotype, such that they also undergo recombination and mutation through the generations, and they evolve alongside the weights $\beta$ and $\gamma$. By attaching these $\sigma$ values, the genotypes become
$Y_{i}=[\beta_{1},\gamma_{1},\beta_{2},\gamma_{2},\cdots,\beta_{p},\gamma_{p},~{}\sigma_{1},\sigma_{2},\cdots,\sigma_{2p}]\,.$ (6)
Each $\sigma$ is normally distributed on $(0,1]$ and mutation occurs as follows:
$\sigma_{i}^{\prime}=\sigma_{i}\,e^{\tau^{\prime}\,N(0,1)+\tau\,N_{i}(0,1)}$ (7)
$\downarrow$
$\beta_{i}^{\prime}=\beta_{i}+\sigma_{i}^{\prime}\,N(0,1)$ (8)
$\gamma_{i}^{\prime}=\gamma_{i}+\sigma_{i+1}^{\prime}\,N(0,1)\,.$ (9)
The variables $\tau$ and $\tau^{\prime}$ are usually set to be inversely proportional to the square root of the population size but can be defined by the user, and can be seen as a learning rate [15]. Furthermore, $\sigma$ is mutated by a log-normal distribution to meet certain requirements, like smaller modifications, the equal likelihood of drawing a value and its reciprocal [32]. After these values evolve we enforce a minimum threshold such that $|\sigma^{\prime}|\geq\sigma_{\rm min}$. If recombination or mutation parameters push $\beta$ and $\gamma$ out of the predefined ranges, we enforce periodic boundaries to return the genes to their proper domain.
After the new generation is created we implement generational survivor selection, where $\mu$ fittest individuals from the older generation will be preserved in the newer generation, if no offspring can surpass their fitness values. This ensures that the best-performing individuals are not lost in evolution and that their genes continue to be passed on through generations. This strategy is particularly useful in noisy environments like quantum hardware, where this safeguards a standard fitness, in such a case even if the population’s average fitness happens to decrease, the next generations can always recuperate and spread the fittest genes again. The operations described are repeated in this order for a fixed number of generations, $g$. Finally, the pseudo-code for our E-QAOA is summarized in the steps of Algorithm 1.
Algorithm 1 : The steps for evolving a population
$Y\leftarrow[Y_{1},Y_{2},\cdots,Y_{n_{\rm pop}}]$
$g\leftarrow 0$
while $\texttt{generations}<g$ do
     for i in (1, $n_{\rm pop}$) do
         $\texttt{fitness${}_{Y_{i}}$}\leftarrow\texttt{f(QAOA($Y_{i}$))}$)
[…]
              $Y_{i}^{\prime}[j]\leftarrow\texttt{mutation($Y_{i}^{\prime}[j],p_{\sigma}$)}$
         end for
     end for
     $Y^{\prime}\leftarrow\texttt{elitism(parents, $Y^{\prime}$, $\mu$)}$
     $g\leftarrow g+1$
end while
The approach for a multi-population EA then follows. Each QPU evolves an independent population of solutions, as we have described above, for a fraction of the total number of generations $g_{\rm f}$, e.g. $g_{\rm f}=g/3$, at which point the fittest individuals from each population migrate to their non-native population. In our case, the fittest individuals are copied over, and they replace the weakest individuals to maintain a constant population size. In this case, the fittest individuals to migrate are chosen by looking at the absolute best Max-Cut given by their distribution, rather than their $\alpha$-tail fitness given by CVaR. The weakest individuals are also evaluated this way before being replaced. Once these switches are done, the independent evolutions continue as normal, with migrations happening every $g_{\rm f}$ generations until termination. These migrations happen classically and their execution is therefore affected by potential queue times and scheduling distributions manually. As mentioned earlier, one of the properties of an MGA approach is to increase diversity within a population. We measure this diversity by looking at uniqueness of values, in particular the number of unique fitness values and unique genes across the population. We normalize this value by the size of the population:
Uniqueness ratio=Number of unique values $\textrm{Uniqueness~{}ratio}=\frac{\text{Number of unique values found}}{n_{\rm pop}}.$ (10)
III Set-up
The simulations for the experiments were done on the Scafell Pike high-performance computing system, hosted by Sci-Tech Daresbury (UK), equipped with dual-socket nodes with a 16-core Intel Xeon Gold CPU per socket. The hardware runs were performed on IBM’s 127-qubit Eagle processors ibm_cusco and ibm_nazca, which have an Error Per Layered Gate (EPLG) for a 100-qubit chain [33] of $5.9\%$ and $3.2\%$ respectively. Within these processors, the QAOA was set to run with $10^{4}$ shots per circuit. For reproducibility, we summarize all the parameters used in Tab. I, and they remain the same for all experiments unless stated otherwise.
TABLE I: List of all variables used, their description and values assigned for the hybrid E-QAOA
Variables Description Value
$n$ Number of nodes (graph size) $[4,\dots,26]$
$n_{\rm pop}$ Number of individuals in a population $[8,\dots,20]$
$g$ Number of generations for termination {10, 20}
$g_{\rm f}$ Number of generations for migration {5, 7}
$p$ Number of rounds for the QAOA circuit 2
$\gamma$, $\beta$ Ansatz parameters to optimize $(-\pi$, $\pi]$
$\alpha$ CVaR value 0.15
$p_{\sigma}$ Probability of mutation 0.2
$\sigma$ Mutation parameter $N(0,1)$
$\sigma_{\rm min}$ Minimum absolute value for $\sigma$ 0.1
$\mu$ Number of elite 1
$\tau$ Parameter used for mutation $\sqrt{2}/2\cdot{n_{\rm pop}}^{-1/4}$
$\tau^{\prime}$ Parameter used for mutation $\sqrt{2}/2\cdot{n_{\rm pop}}^{-1/2}$
IV Results and Discussion
Now, we present the results of our experiments and a discussion on their performance. Tests were first run on a simulator to quickly evaluate potential advantages and tune the appropriate parameters. Then, they were run on quantum hardware.
IV-A Simulations
In Fig. 2 we show the performance of our EA optimizer against COBYLA, for increasing graph sizes. On the top plot, we note that for small graphs both methods achieve the optimal solution, i.e. an approximation ratio of 1. Whereas from $n\geq 16$, both optimizers start decreasing in accuracy, the EA maintains an advantage over COBYLA for both accuracy and variance over an average of 10 runs. In this plot CVaR was used which the reader can note is significantly more stable than using the max_count option. The latter is shown in the bottom plot, where the approximation ratio reaches a value of 1 fewer times, and in general, this method is shown to be less stable. For example, for $n=12$ a particularly hard-to-solve graph appears to have been randomly chosen. Conversely, both optimizers solved graphs with the following two sampled sizes more efficiently. In this plot, the variance for COBYLA is substantially increased and the accuracy takes a large hit, dropping as low as  $\approx\,0.61$ for both $n=20$ and $26$. On the other hand, the EA optimizer is shown to be more robust, even though the accuracy does not remain as high as with CVaR, the variance increases only slightly, and it is still able to maintain an approximation ratio at  $\approx\,0.77$ or above for all cases, with a much lower discrepancy between repetitions than COBYLA. Independently of the fitness function used the EA optimizer always performs on par or better than COBYLA and with a smaller variance
Figure 2: Approximation ratio (3), averaged over 10 runs, for graph sizes $n=[4,10,12,14,16,20,26]$ on $d$-3 regular graphs solved using QAOA paired with either COBYLA or the EA. For sub-figure a) CVaR is used, whilst b) uses max_count to calculate the fitness of solutions. We used $g=10$, $n_{\rm pop}=10$ for all simulations, and the standard deviation for error bars. The horizontal dotted red line represents the maximum approximation ratio achievable.
We theorize this higher robustness for the EA is due to an evolving population of solutions rather than just an individual one, and the fact that the EA is less susceptible to initial conditions, e.g., due to a poor initial guess. Furthermore, the EA can explore a wider search space within the fitness landscape instead of being guided by gradients and is thus able to consistently find more accurate solutions.
Figure 3: Simulated approximation ratio as a function of the population size ($n_{\rm pop}$) for 20-node (left) and 26-node (right) graphs. Simulations were run for values of $n_{\rm pop}=[8,10,12,16,20]$, indicated by the vertical dotted lines with labels. For a given $n_{\rm pop}$, the approximation ratio for single-population runs (black) is compared to multi-population-EA results for $g_{\rm f}=5$ (red) and $g_{\rm f}=7$ (green). Markers represent the mean approximation ratio from 10 independent regular $d$-3 random graphs and the error bars span the minimum and maximum obtained for each $n_{\rm pop}$. The black and green markers are slightly shifted along the $x$-axis to facilitate the comparison; as guidelines, we show solid horizontal lines (and grey boxes) next to the black markers to compare the mean multi-population ratios (and the extent) with the single-population results. The hatched area for ratios above 1 is by definition excluded.
In Fig. 3 we show the simulations for finding the average Max-Cut approximation ratio on a graph with $n=20$ nodes on the left, and $n=26$ on the right, for a single population algorithm versus a multi-population one as explained in the methodology. Note that the population size is the same for each isolated population, thus the MGA approach is effectively evolving double the number of individuals independently. In the left plot, the multi-population algorithm for $g_{\rm f}=5$, shown in red, achieves a slightly better solution quality than the single population for all combinations of population size, over an average of 10 random graphs. For both the single and multi-strategy the mean approximation ratio tends to increase as the population size increases, except for $n_{\rm pop}=16$. This is expected as evaluating more individuals and increasing the population size, should help, but is not necessarily a linear relationship and is considered a tunable parameter. For $g_{\rm f}=7$ this advantage slightly increases, over the one achieved with $g_{\rm f}=5$, in four out of the 5 instances. Beyond achieving a better mean solution, the multi-population strategies also achieve a higher minimum for all experiments. Another interesting thing to observe is that the multi-population algorithm, for both $g_{\rm f}$ values, will in most cases shown, achieve at least the same accuracy rate as the single-population algorithm with a larger population size. This is except for $n_{\rm pop}=16$ again, yet, in the case of $n_{\rm pop}=12$ the multi-population approach with $g_{\rm f}=5$ achieves an accuracy better than the single population one with $n_{\rm pop}=20$. This result suggests there is potential for a speed-up for the MGA approach, but such an investigation of that performance is outside the scope of this paper.
On the right plot instead, the results and advantages are not as simple to interpret. In this case, for $n=26$, there needs to be some tuning between $n_{\rm pop}$ and $g_{\rm f}$ to achieve the best approximation ratio. For the multi-population with $g_{\rm f}=5$, we still achieve better or on-par results, with a lower population size than the single population does with $n_{\rm pop}=20$. A larger $g_{\rm f}$ value only helps with larger population sizes, otherwise, it is detrimental. In these cases, for $n_{\rm pop}\leq 12$, waiting for a larger number of generations might hinder successful evolution and convergence as we are introducing a new fresh individual when the algorithm has already started converging towards another set of genes. This competition between two very fit individuals might split a low-size population in a way that does not explore the fitness landscape optimally. In both plots, the multi-population algorithm with $g_{\rm f}=7$ and $n_{\rm pop}=20$ achieves the highest average approximation ratio overall.
IV-B Hardware
Following the simulation results, we performed an initial experiment on the ibm_nazca device, solving the Max-Cut problem on 5 random $d$-3 regular graphs of size $n=16$. Consistent with simulation, the optimal Max-Cut was found in all cases. Given this, we scaled up the experiments to graphs of size $n=20$ with a population size of $n_{\rm pop}=12$. This test case was chosen because it gave some of the most promising simulation results for the multi-population strategy. As can be seen in Fig. 4, we tested the single and multi-population approach with $g_{\rm f}=5$. Due to constraints on QPU time, we ran these on the hardware for 15 generations rather than 20, and tested on 5 random graphs rather than 10. While this limits any claims we can make regarding the advantages of E-QAOA and multi-population strategies, we present the hardware results nonetheless because they serve as a proof-of-concept and validation of our simulations. The multi-population strategy achieves a slightly higher mean and a higher maximum on hardware than the single population. Both methods perform slightly worse on hardware than in simulation. However, this is expected and it is promising to see that the gap between hardware and simulation is relatively small, and that the small improvement of MGA over single population is maintained on hardware.
Figure 4: As in Fig. 3, but for hardware runs (empty markers) of 5 random regular $d$-3 graphs with $n=20$. Hardware runs configured with $n_{\rm pop}=12$, $g=15$ and $g_{\rm f}=5$. The single-population run (empty square) was performed on ibm_nazca and the multi-population run (empty triangle) was performed on both ibm_nazca and ibm_cusco. For comparison, we show simulation results for $n_{\rm pop}=12$ and 20 as filled markers.
Figure 5: Both figures show a uniqueness ratio (10) over the evolution of generations for a single-population (solid red) and a multi-population (dashed blue) on the average of 5 random graphs of size $n=20$. For all hardware runs $g=15$, $g_{\rm f}=5$ and $n_{\rm pop}=12$. These statistics are achieved on the ibm_nazca machine, with the shaded area representing the standard deviation. Sub-figure a) shows the uniqueness in fitness values calculated with CVaR. Sub-figure b) shows the uniqueness in the first allele of the genetic code, i.e. $\beta_{0}$. Note that for the multi-population approach, the uniqueness is calculated only on one population, within the ibm_nazca machine, as such the values between the two approaches overlap before the first migration. The vertical dotted lines represent the generation number at which these migrations happen. Finally, we use a spline function to smooth the uniqueness ratio values between (discrete) generations.
As mentioned in the introduction a multi-population approach serves to amplify the diversity within the genes of a population, avoid premature convergence, and search the fitness landscape for values that might not otherwise be reached with a single population. Fig. 5 shows how the uniqueness for two different parameters evolves for the single and multi-population approaches as tested on hardware. On the left side, sub-figure (a) shows the uniqueness with respect to how many fitness values are found within the ibm_nazca population. In this instance, the multi-population approach trends slightly higher whilst inter-crossing with the single-population approach. Given that the fitness value is not related in a one-to-one fashion to the genes or ansatz parameters of the circuit, the fitness value cannot show the whole picture on its own, since a family of $\beta$ and $\gamma$ combinations would produce the same CVaR output when evaluating the fitness. To provide more complete information, we also show the uniqueness of $\beta_{0}$ values in the right sub-figure (b). In this case, the separation is more pronounced, with the migrations, shown in the vertical dotted lines happening at regular intervals, pushing and maintaining a higher uniqueness ratio within the multi-population approach. The comparison is done using only one allele as the cross-over operator for recombination, which is happening on all alleles, should generate unique individuals in these terms without the aid of mutation. As for the uniqueness of fitness values, diversity between genes does not provide the whole picture either, and would require further analysis of the expressivity of the circuit for the genes, and a fitness value using only the best Max-Cut value within the distribution, i.e. $\alpha=10^{-4}$, or 1 count in $10^{4}$ shots.
Given these results, we can conclude that the multi-population E-QAOA has potential but may need careful parameter tuning to produce optimal results and successfully compete with the single population E-QAOA strategy. In this respect, we would allow further generations to evolve and converge without migrations once successful swaps have already modified the distributed populations. This further step is necessary to ultimately allow diversity within the population to fall, and let individual solutions converge towards the best genetic codes.
V Conclusion
In this paper, we presented results comparing an evolutionary approach against a standard QAOA using COBYLA, on a simulator for $d$-3 regular graphs, ranging in size from 4 to 26 nodes. We studied the accuracy and variance for the approximation ratio, using both CVaR and max_count, accounting for the former being a more stable fitness evaluator. Next, we noted the best EA parameters to take forward in producing a multi-population E-QAOA, testing this MGA approach both on the simulator and hardware. On the simulator, we tested two different migration schedules to see how these would affect performance and then tested the multi-population E-QAOA with $g_{\rm f}=5$ on the hardware. Results between simulators and hardware were in accordance, even though, as expected, hardware results produced poorer accuracy rates. Lastly, we investigated the uniqueness of fitness and genetic code within a population using a single- and multi-population approach to guarantee an increase in genetic diversity in the latter strategy.
Our experiments, although performed on small-size problems with limited computational resources, have paved the way for future work. Despite these limitations, our work presents a novel approach by using independent isolated populations to evolve E-QAOA and distributing them on 2 scheduled QPUs by running experiments on quantum hardware. Further investigations of multi-population E-QAOA approaches are already underway. These upcoming results aim to consolidate the novel techniques presented in this paper, characterize their advantages, and further probe parallelization techniques. We have set the foundations for this work to eventually produce advancements in speed-ups alongside also boosting accuracy.
Acknowledgment
This work was supported by the Hartree National Centre for Digital Innovation, a UK Government-funded collaboration between STFC and IBM. IBM, the IBM logo, and www.ibm.com are trademarks of International Business Machines Corp., registered in many jurisdictions worldwide. Other product and service names might be trademarks of IBM or other companies. The current list of IBM trademarks is available at www.ibm.com/legal/copytrade.
The research in this paper made use of the following software packages and libraries: Python [34], Qiskit [27], Numpy [35], Scipy [28], Random [36], NetworkX [37, 38], Matplotlib [39, 40].